OpenStudy (anonymous):

FAN AND MEDAL*** (HELP) 4)how do you find the zeros of a function by graphing? or, where are the zeros (solutions) of a function located on the graph of the function. 5) which do you prefer long division or synthetic division? why?

OpenStudy (anonymous):

this question does not make any sence to me, sorry

OpenStudy (anonymous):

ik ;(

hero (hero):

In other words, the zeroes of a polynomial function occur on which axis?

OpenStudy (anonymous):

idontknow @hero ;/

hero (hero):

The zeroes of a function is the x-value that makes the function equal to zero. For example: Suppose you have f(x) = 3x^2 + 6x. In order to find the zero for this, we would (starting with the vertex), continue plotting plotting points along the graph on either side of the x-axis and then approximate its value based on the graph. Then you'd keep approximating until you found the actual zero.

hero (hero):

That's the graphing technique, however, its much easier to simply factor 3x^2 + 6x, set it equal to zero, and then use the product property rule to find the zeroes. In this case, 3x^2 + 6x = 0 3x(x + 2) = 0 3x = 0 x + 2 = 0 x = 0 x = -2 So 0 and -2 are the zeroes of the function. Graphically, these zeroes occur on the x-axis.

hero (hero):

When thinking about the zeroes of a function, the question you want to ask yourself is, "What is the value of x when f(x) equals zero?".